# Personalized Generation In Large Model Era: A Survey

### v1

@@ 3.6.4 Evaluation Metrics
To quantify personalization in text generation tasks such as personal assistant and comment generation, existing studies commonly employ (1) term-matching metrics, such as ROUGE, BLEU, Meteor, CIDEr Cohen et al. (2022); Alaluf et al. (2025); Nguyen et al. (2024b); Pi et al. (2024); An et al. (2024); (2) semantic matching metrics like CLIPScore Cohen et al. (2022); (3) recall, precision and F1-score to validate whether the user-specific concept appears in the generated caption Cohen et al. (2022); Alaluf et al. (2025); Nguyen et al. (2024b); Pi et al. (2024); An et al. (2024); (4) human evaluation to determine alignment with ground truths in terms of emotion, style, and relevance.
To measure whether agents align with diverse human preferences, studies employ success rate, success weighted by path length, distance to goal, and episode length Poddar et al. (2024); Hwang et al. (2024). Human evaluation is also an effective validation method for determining whether an agent has successfully completed a personalized task Hwang et al. (2024); Cui et al. (2024b).
@@ 5.1 Technical Challenges
  • Deliberative Reasoning for PGen. In certain personalized scenarios that prioritize content quality over temporal efficiency – such as digital advertising, where advertisers typically serve only several ads per user each day – inference scaling presents significant opportunities for enhancing user satisfaction. Prior work mainly focuses on multi-turn refinement to progressively enhance content relevance and personalization Nabati et al. (2024). Inspired by the great success of LLM reasoning Guan et al. (2024); Guo et al. (2025), deliberative reasoning may emphasize extensive logical and contextual reasoning beforehand, enabling a thorough analysis of user preferences to drive more effective content personalization Fang et al. (2025).
@@
  • Multi-modal Personalization. Existing PGen research mainly focuses on single-modality generation, while multi-modal personalization remains underexplored, such as personalized social media posts that integrate both image and text. This challenge requires high-quality, instruction-aligned, and personalized output while ensuring consistency across multiple modalities.
@@ 5.3 Trustworthiness
  • Privacy. PGen relies on user-specific data for content personalization, raising concerns about privacy issues. Striking a balance between effective personalization and robust privacy protection is crucial for advancing this field.
  • Fairness and Bias. PGen may unintentionally reinforce biases and stereotypes present in training data, leading to skewed or discriminatory outcomes. Effective detection and mitigation strategies are essential to protect diverse user groups.

### v2

@@ 3.7.4 Evaluation Metrics
To quantify the degree of personalization in text generation tasks such as personal assistant and comment generation, existing studies commonly employ (1) term-matching metrics, such as ROUGE, BLEU, Meteor, CIDEr Cohen et al. (2022); Alaluf et al. (2025); Nguyen et al. (2024b); Pi et al. (2024); An et al. (2024); (2) semantic matching metrics, such as CLIPScore Cohen et al. (2022); (3) recall, precision and F1-score to validate whether the user-specific concept appears in the generated caption Cohen et al. (2022); Alaluf et al. (2025); Nguyen et al. (2024b); Pi et al. (2024); An et al. (2024); (4) human evaluation to determine alignment with ground truths in terms of emotion, style, and relevance.
To measure whether agents align with diverse human preferences, studies employ success rate, success weighted by path length (SPL), distance to goal, and episode length Poddar et al. (2024); Hwang et al. (2024). Human evaluation is also an effective validation method for determining whether an agent has successfully completed a personalized task Hwang et al. (2024); Cui et al. (2024b).
@@ 5.1 Technical Challenges
  • Deliberative Reasoning for PGen. In certain personalized scenarios that prioritize content quality over temporal efficiency – such as digital advertising, where advertisers typically serve only several ads per user each day – inference scaling presents significant opportunities for enhancing user satisfaction. Prior work mainly focuses on multi-turn refinement to progressively enhance content relevance and personalization Nabati et al. (2024). Inspired by the great success of LLM reasoning Guan et al. (2024); Guo et al. (2025); Peng et al. (2025), deliberative reasoning may emphasize extensive logical and contextual reasoning beforehand, enabling a thorough analysis of user preferences to drive more effective content personalization Fang et al. (2025).
@@
  • Mitigating Filter Bubbles. PGen may inadvertently reinforce users’ existing preferences or beliefs and contribute to polarization Lazovich (2023), which has been widely studied in recommender systems. Strategies such as diversity enhancement Chen et al. (2018a) and user-controllable inference Wang et al. (2022b) offer valuable insights for mitigating this in PGen. Moreover, multi-agent generation introduces diverse perspectives through distinct generative personas, providing a promising solution within generative settings Zhang et al. (2024j).
  • User Data Management. As the cornerstone of PGen, user data management is essential for collecting, storing, and structuring user-specific data, as well as enabling user-controllable personalization. Existing approaches primarily adopt either an external user memory module combined with RAG to manage user data Mysore et al. (2024) or embed user-specific information directly into model parameters to achieve personalization Ruiz et al. (2023). However, developing a more effective and efficient strategy for lifelong user data management remains an open problem.
  • Multi-modal Personalization. Existing PGen research mainly focuses on single-modality generation, while multi-modal personalization remains underexplored, such as personalized social media posts that integrate both image and text. This challenge requires high-quality, instruction-aligned, and personalized output while ensuring consistency across multiple modalities. Recently, the emergence of unified multimodal models, such as Show-o Xie et al. (2025), Transfusion Zhou et al. (2025a), Janus-Pro Chen et al. (2025), BAGEL Deng et al. (2025), and GPT-4o OpenAI (2024), which provide shared architectures for personalized multimodal generation and open opportunities for transferring personalization techniques across modalities by leveraging insights from single-modality methods.
@@ 5.3 Trustworthiness
  • Privacy. PGen relies on user-specific data for content personalization, raising concerns about privacy issues. Striking a balance between effective personalization and robust privacy protection is crucial for advancing this field. Prior approaches, such as on-device personalization Wang and Chau (2024); Cho et al. (2024); Bang et al. (2024), federated learning Jiang et al. (2024a); Zhang et al. (2024h), differential privacy Flemings et al. (2024), and adversarial perturbation Van Le et al. (2023), offer promising strategies for achieving this balance, enabling personalization while safeguarding user privacy.
  • Fairness and Bias. PGen may unintentionally reinforce biases and stereotypes present in training data, leading to skewed or discriminatory outcomes Wan et al. (2023). Extensive studies have explored debiasing strategies for generative models, including context steering Pandey et al. (2024), structured prompt Furniturewala et al. (2024); Gallegos et al. (2025), and causal-guided biased sample identification Sun et al. (2024b), which offer valuable insights for mitigating potential biases in PGen.
@@ 6 Conclusion
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China (62272437 and U24B20180), in part by Center for Intelligent Information Retrieval, in part by NSF grant number 2143434, and in part by an award from Google. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsor.